# Robust Distributed Estimation: Extending Gossip Algorithms to Ranking and Trimmed Means

### v6

@@ 3.2 Numerical Experiments
(a) Role of $\sigma_{n}(\cdot)$
@@
Figure 2: Illustration of the behavior of GoRank: (a) shows how the absolute error of the rank estimates of GoRank aligns with the shape of the function $x\mapsto\sqrt{x(1-x)}$, where $x=(r-1)/n$ for a given rank $r$, and highlights the role of the dispersion measure $\sigma_{n}$ in the error bound; (b) compares the convergence rate of GoRank across different graph topologies (i.e., levels of connectivity), illustrating the influence of the constant $c$ in the bound; (c) compares GoRank with existing method (Baseline) and an alternative ranking algorithm developed in this work (Baseline++).
Setup.
We conduct experiments on a dataset $S=\{1,\dots,n\}$ with $n=500$, distributed across nodes of a communication graph. Our evaluation metric is the normalized absolute error between estimated and true ranks, i.e., for node $k$ at iteration $t$, the error is defined as $\ell_{k}(t)=|R_{k}(t)-r_{k}|/n$. While more sophisticated ranking metrics such as Kendall $\tau$ exist, they are not relevant in our context, as we focus on individual rank estimation accuracy. We first examine the impact of the constants appearing in our theoretical bounds (see Theorem 1), the one related to graph connectivity and the one reflecting rank centrality. To this end, we consider three graph topologies: the complete graph ($c=4.01\times 10^{-3}$), in which every node is directly connected to all others, yielding maximal connectivity; the two-dimensional grid ($c=1.65\times 10^{-5}$), where each node connects to its four immediate neighbors; and the Watts–Strogatz network ($c=3.31\times 10^{-4}$), a randomized graph with average degree $k=4$ and rewiring probability $p=0.2$, offering intermediate connectivity between the complete and grid graphs. Then, we compare GoRank with the two other ranking algorithms Baseline and Baseline++. Figure (a) was generated on a Watts-Strogatz graph, averaged over $31e$ 3 trials. Figure (b) and (c) show the convergence of ranking algorithms with mean and standard deviation computed from $100$ trials: figure (b) shows the convergence of GoRank on different graph topologies; figure (c) compares the convergence of the ranking algorithms on a 2D Grid graph. All experiments were run on a single CPU with 32 GB of memory for $48e$ 4 iterations, with a total execution time of approximately two hours. The code for our experiments will be publicly available. [^1] The code is available in the Supplementary Material.
@@ 4.1 Algorithm – Convergence Analysis
5:        Update rank: $R_{k}\leftarrow\texttt{rank.update}(k,t)$.
@@
We introduce GoTrim, a gossip algorithm to estimate trimmed mean statistics (see Algorithm 2), which dynamically computes a weighted average using current estimated ranks. Notably, GoTrim can be paired with any ranking algorithm, including GoRank consequently. Let $Z_{k}(t)$ and $W_{k}(t)$ denote the local estimates of the statistic and weight at node $k$ and iteration $t$. First, by Equation 2, the $\alpha$-trimmed mean can be computed via standard gossip averaging: for all $k\in[n]$, $Z_{k}(t)=(1/n)\sum_{l=1}^{n}W_{l}(t)\cdot X_{l}$, where $W_{l}(t)=w_{n,\alpha}(R_{l}(t))$ with $w_{n,\alpha}(\cdot)$ defined in Section 2. Secondly, since ranks $R_{k}(t)$ vary over iterations, the algorithm dynamically adjusts to correct past errors: at each step, it compensates by injecting $\left(W_{k}(t)-W_{k}(t-1)\right)\cdot X_{k}$ into the averaging process. On the one hand, we have an averaging operation which is captured by the averaging matrix: $\mathbf{W}_{2}(t)=\mathbf{I}_{n}-(e_{i}-e_{j})(e_{i}-e_{j})^{\top}/2$. Similarly to the permutation matrix (see the previous section), the expectation of the averaging matrix is symmetric, doubly stochastic and has spectral gap that satisfies $10<c_{2}<$ 1 and is given by $c_{2}=c/2$ [31, 3]. On the other hand, we have a non-linear operation that depends on the estimated ranks: at each iteration $t>0$, each node $k$ is updated as $Z_{k}(t)=Z_{k}(t-1)+\delta_{k}(t)\cdot X_{k}$, where $\delta_{k}(t)=W_{k}(t)-W_{k}(t-1)$. Hence, the evolution of the estimates can be expressed as $\mathbf{Z}(t)=\mathbf{W}_{2}(t)\left(\mathbf{Z}(t-1)+\boldsymbol{\delta}(t)\odot\boldsymbol{X}\right),$ where $\mathbf{Z}(t)=(Z_{1}(t),\ldots,Z_{n}(t))$ and $\boldsymbol{\delta}(t)=(\delta_{1}(t),\ldots,\delta_{n}(t))$. Taking the expectation over the sampling process, the expected estimates are given by: $\mathbb{E}[\mathbf{Z}(t)]=\mathbf{W}_{2}\left(\mathbb{E}[\mathbf{Z}(t-1)]+\Delta\boldsymbol{w}(t)\odot\boldsymbol{X}\right)$ with $\mathbf{W}_{2}=\mathbf{I}_{n}-(1/2|E|)\mathbf{L}$ and $\Delta\boldsymbol{w}(t)=\mathbb{E}[\boldsymbol{\delta}(t)]$. For $t=1$, since $Z_{k}(0)=0$, we have $\mathbb{E}[\boldsymbol{Z}(1)]=\mathbf{W}_{2}\Delta\boldsymbol{w}(1)\odot\boldsymbol{X}$. Recursively, for any $t>0$,
@@
Let $\mathbf{R}(t)$ and $\mathbf{W}(t)$ be defined as in Algorithm 1 and Algorithm 2, respectively. For all $k\in[n]$ and $t>0$, we have:
$\left|\mathbb{E}[W_{k}(t)]-w_{n,\alpha}(r_{k})\right|\leq\mathcal{O}\left(\frac{1}{\gamma_{k}^{2}{ct}}\right)\cdot\sigma_{n}(r_{k})^{2},$
where $\gamma_{k}=\min(\left|r_{k}-a\right|,\left|r_{k}-b\right|)\geq 1/2$ with $a=\lfloor\alpha n\rfloor+1/2$ and $b=n-\lfloor\alpha n\rfloor+1/2$ being the endpoints of interval $I_{n,\alpha}$. The constants $c$ and $\sigma_{n}(\cdot)$ are those defined in Theorem 2.
@@
Let $\mathbf{Z}(t)$ be defined as in Algorithm 2, and assume the ranking algorithm is Algorithm 1. Then, for any $t>T^{*}=\min\left\{t>1\,\middle|\,ct>2\log(t)\right\}$,
@@
where $\mathbf{K}=(\sigma_{n}(r_{1})^{2}/\gamma_{1}^{2},\;\ldots,\;\sigma_{n}(r_{n})^{2}/\gamma_{n}^{2})$ and $c$, $\sigma_{n}(r_{k})$, $\gamma_{k}$ are the constants defined in Lemma 1.
@@ 4.2 Robustness Analysis - Breakdown Points
Consider a dataset $S$ of size $n\geq 1$ and $T_{n}=T_{n}(S)$ a real-valued statistic based on it. The breakdown point for the statistic $T_{n}(S)$ is defined as $\varepsilon^{*}=p_{\infty}/n$, where $p_{\infty}=\min\{p\in\mathbb{N^{*}}:\,\sup_{S^{\prime}_{p}}|T_{n}(S)-T_{n}(S^{\prime}_{p})|=\infty\}$, the supremum being taken over all corrupted datasets $S^{\prime}_{p}$ obtained by replacing $p$ samples in $S$ by arbitrary samples. The breakdown point is a popular notion of robustness [19], corresponding here to the fraction of samples that need to be corrupted to make the statistic $T_{n}$ arbitrarily large (i.e., "break down"). For example, the breakdown point of the $\alpha$-trimmed mean is given by $\lfloor\alpha n\rfloor/n$, which is approximately $\alpha$. We consider here a generalization of this notion, namely the $\tau$-breakdown point by replacing $p_{\infty}$ with $p_{\tau}=\min\{p\in\mathbb{N^{*}}\enspace\text{, }\sup_{S^{\prime}_{p}}|T_{n}(S)-T_{n}(S^{\prime}_{p})|\geq\tau\}$, where $\tau>0$ is a threshold parameter. Since our algorithm does not compute the exact $\alpha$-trimmed mean, we focus instead on determining the $\tau$-breakdown point $\varepsilon^{*}_{k}(t)$ of the partial $\alpha$-trimmed mean at iteration $t$ for each node $k$, i.e., the estimate $Z_{k}(t)$. Given that this quantity was previously shown to converge to the $\alpha$-trimmed mean, we expect that $\varepsilon^{*}_{k}(t)\leq\alpha$. The following theorem provides framing bounds for the breakdown point of the estimates of GoTrim when paired with GoRank.
@@
See Section E of the Supplementary Material for the technical proof, which relies on the idea that, when estimating a partial $\alpha$-trimmed mean, there are two sources of error: (1) the uncertainty in rank estimation, which can lead to incorrect data points being included in the mean, and (2) the delay from the gossip averaging, which requires a certain propagation time $T$ to update the network estimates. Note that, as $t\rightarrow\infty$, Theorem 4 recovers the breakdown point of the exact $\alpha$-trimmed mean.
@@ 4.3 Numerical Experiments
(a) Role of $\sigma_{n}(r_{k})^{2}/\gamma_{k}^{2}$
@@ Results.
5 Conclusion and Future Work
We introduced and analyzed two novel gossip algorithms: GoRank for rank estimation and GoTrim for trimmed mean estimation. We proved convergence rates of $\mathcal{O}(1/t)$ and established robustness guarantees for GoTrim through breakdown point analysis. Empirical results show both methods perform well on large, poorly connected networks: GoRank quickly estimates ranks, and GoTrim is robust to outliers and improves on the naive mean. Although our analysis focuses on the synchronous setting, real-world systems are often asynchronous. We present the asynchronous version of GoRank in the Supplementary Material; however, similar to Baseline++, its convergence analysis is left for future work. In addition, extending GoTrim for decentralized estimation of rank statistics [20], key tools in data analysis, e.g., for robust hypothesis testing, is another promising direction. See the Supplementary Material for further discussion.
@@ Appendix A Gossip Algorithms for Ranking
Here, we detail two gossip algorithms for ranking, which can be considered natural competitors to GoRank. The algorithm presented first was proposed in [7] and, to our knowledge, is the only decentralized ranking algorithm documented in the literature. The algorithm presented next can be seen as a variant of the latter, incorporating the more efficient communication scheme of GoRank.
@@
Chiuso et al. propose a gossip algorithm for distributed ranking in a general (connected) network [7]. They demonstrate that this algorithm solves the ranking problem almost surely in finite time. However, they do not provide any convergence rate or non-asymptotic convergence results. The algorithm is outlined in Algorithm 3 and proceeds as follows. At each time step $t$, an edge $(i,j)$ is selected, and the algorithm operates in three phases:
@@ B.1 Properties of Gossip Matrices
    The eigenvalue $\lambda_{2}(\alpha)$ satisfies $10\leq\lambda_{2}(\alpha)<$ 1 and $\lambda_{2}(\alpha)=1-\frac{\lambda_{2}}{\alpha|E|}$ where $\lambda_{2}$ the spectral gap (or second smallest eigenvalue) of the Laplacian.
@@
The proofs of (a), (b), (d), and (e) are omitted for brevity. For more details, we refer the reader to [3, 8]. The proof of (c) follows from the fact that
@@
The proof of (f) can be found in [8] (see Lemmas 1, 2 and 3). ∎
@@ B.2 Convergence Analysis of Gossip for Averaging
Algorithm 5 Gossip algorithm for estimating the standard average [3]
@@
Let us assume that $\mathcal{G}=(V,E)$ is connected and non bipartite. Then, for $\mathbf{Z}(t)$ defined in Algorithm 5, we have that for all $k\in[n]$ :
@@
The proof is kept brief, as this is a standard result. For more details, we refer the reader to [3, 8]. To prove this lemma, we will primarily use points (e) and (f) from Lemma.
@@
Setting $c_{2}=1-\lambda_{2}(2)>0$ finishes the proof, as $10<\lambda_{2}(2)<$ 1. ∎
@@
Let us assume that $\mathcal{G}=(V,E)$ is connected and non bipartite. Then, for $\mathbf{Z}(t)$ defined in Algorithm 5, we have that for any $t>0$ :
@@
The original proof can be found in [3].
@@
Setting $c_{2}=1-\lambda_{2}(2)>0$ finishes the proof, as $10<\lambda_{2}(2)<$ 1. ∎

### v7

@@ 3.2 Numerical Experiments
Setup.
We conduct experiments on a dataset $S=\{1,\dots,n\}$ with $n=500$, distributed across nodes of a communication graph. Our evaluation metric is the normalized absolute error between estimated and true ranks, i.e., for node $k$ at iteration $t$, the error is defined as $\ell_{k}(t)=|R_{k}(t)-r_{k}|/n$. While more sophisticated ranking metrics such as Kendall $\tau$ exist, they are not relevant in our context, as we focus on individual rank estimation accuracy. We first examine the impact of the constants appearing in our theoretical bounds (see Theorem 1), the one related to graph connectivity and the one reflecting rank centrality. To this end, we consider three graph topologies: the complete graph ($c=4.01\times 10^{-3}$), in which every node is directly connected to all others, yielding maximal connectivity; the two-dimensional grid ($c=1.65\times 10^{-5}$), where each node connects to its four immediate neighbors; and the Watts–Strogatz network ($c=3.31\times 10^{-4}$), a randomized graph with average degree $k=4$ and rewiring probability $p=0.2$, offering intermediate connectivity between the complete and grid graphs. Then, we compare GoRank with the two other ranking algorithms Baseline and Baseline++. Figure (a) was generated on a Watts-Strogatz graph, averaged over $1e3$ trials. Figure (b) and (c) show the convergence of ranking algorithms with mean and standard deviation computed from $100$ trials: figure (b) shows the convergence of GoRank on different graph topologies; figure (c) compares the convergence of the ranking algorithms on a 2D Grid graph. All experiments were run on a single CPU with 32 GB of memory for $8e4$ iterations, with a total execution time of approximately two hours. The code for our experiments will be publicly available. [^1] The code is available in the Supplementary Material.
(a) Role of $\phi$
@@
Figure 2: Illustration of the behavior of GoRank: (a) shows how the absolute error of the rank estimates of GoRank aligns with the shape of the function $\phi$, and highlights the role of the constant $\sigma_{k}$ for $k\in[n]$ in the error bound; (b) compares the convergence rate of GoRank across different graph topologies (i.e., levels of connectivity), illustrating the influence of the constant $c$ in the bound; (c) compares GoRank with existing method (Baseline) and an alternative ranking algorithm developed in this work (Baseline++).
@@ 4.1 Algorithm – Convergence Analysis
5:   Update rank: $R_{k}\leftarrow\texttt{rank.update}(k,s)$.
@@
We introduce GoTrim, a gossip algorithm to estimate trimmed mean statistics (see Algorithm˜2), which dynamically computes a weighted average using current estimated ranks. Notably, GoTrim can be paired with any ranking algorithm, including GoRank consequently. Let $Z_{k}(t)$ and $W_{k}(t)$ denote the local estimates of the statistic and weight at node $k$ and iteration $t$. First, by Equation˜2, the $\alpha$-trimmed mean can be computed via standard gossip averaging: for all $k\in[n]$, $Z_{k}(t)=(1/n)\sum_{l=1}^{n}W_{l}(t)\cdot X_{l}$, where $W_{l}(t)=w_{n,\alpha}(R_{l}(t))$ with $w_{n,\alpha}(\cdot)$ defined in Section 2. Secondly, since ranks $R_{k}(t)$ vary over iterations, the algorithm dynamically adjusts to correct past errors: at each step, it compensates by injecting $\left(W_{k}(t)-W_{k}(t-1)\right)\cdot X_{k}$ into the averaging process. On the one hand, we have an averaging operation which is captured by the averaging matrix: $\mathbf{W}_{2}(t)=\mathbf{I}_{n}-(e_{i}-e_{j})(e_{i}-e_{j})^{\top}/2$. Similarly to the permutation matrix (see the previous section), the expectation of the averaging matrix is symmetric, doubly stochastic and has spectral gap that satisfies $0<c_{2}<1$ and is given by $c_{2}=c/2$ Shah et al. [2009], Boyd et al. [2006]. On the other hand, we have a non-linear operation that depends on the estimated ranks: at each iteration $t>0$, each node $k$ is updated as $Z_{k}(t)=Z_{k}(t-1)+\delta_{k}(t)\cdot X_{k}$, where $\delta_{k}(t)=W_{k}(t)-W_{k}(t-1)$. Hence, the evolution of the estimates can be expressed as $\mathbf{Z}(t)=\mathbf{W}_{2}(t)\left(\mathbf{Z}(t-1)+\boldsymbol{\delta}(t)\odot\boldsymbol{X}\right),$ where $\mathbf{Z}(t)=(Z_{1}(t),\ldots,Z_{n}(t))$ and $\boldsymbol{\delta}(t)=(\delta_{1}(t),\ldots,\delta_{n}(t))$. Taking the expectation over the sampling process, the expected estimates are given by: $\mathbb{E}[\mathbf{Z}(t)]=\mathbf{W}_{2}\left(\mathbb{E}[\mathbf{Z}(t-1)]+\Delta\boldsymbol{w}(t)\odot\boldsymbol{X}\right)$ with $\mathbf{W}_{2}=\mathbf{I}_{n}-(1/2|E|)\mathbf{L}$ and $\Delta\boldsymbol{w}(t)=\mathbb{E}[\boldsymbol{\delta}(t)]$. For $t=1$, since $Z_{k}(0)=0$, we have $\mathbb{E}[\boldsymbol{Z}(1)]=\mathbf{W}_{2}\Delta\boldsymbol{w}(1)\odot\boldsymbol{X}$. Recursively, for any $t>0$,
@@
Let $\mathbf{R}(t)$ and $\mathbf{W}(t)$ be defined as in Algorithm˜1 and Algorithm˜2, respectively. For all $k\in[n]$ and $t>0$, we have:
$\left|\mathbb{E}[W_{k}(t)]-w_{n,\alpha}(r_{k})\right|\leq\mathcal{O}\left(\frac{1}{\gamma_{k}^{2}{ct}}\right)\cdot\sigma_{k}^{2},$
where $\gamma_{k}=\min(\left|r_{k}-a\right|,\left|r_{k}-b\right|)\geq 1/2$ with $a=\lfloor\alpha n\rfloor+1/2$ and $b=n-\lfloor\alpha n\rfloor+1/2$ being the endpoints of interval $I_{n,\alpha}$. The constants $c$ and $\sigma_{k}$ are those defined in Theorem 2.
@@
Let $\mathbf{Z}(t)$ be defined as in Algorithm˜2, and assume the ranking algorithm is Algorithm˜1. Then, for any $t>T^{*}=\min\left\{t>1\,\middle|\,ct>2\log(t)\right\}$,
@@
where $\mathbf{K}=(\sigma_{1}^{2}/\gamma_{1}^{2},\;\ldots,\;\sigma_{n}^{2}/\gamma_{n}^{2})$ and $c$, $\sigma_{k}$, $\gamma_{k}$ are the constants defined in Lemma˜1.
@@ 4.2 Robustness Analysis - Breakdown Points
Consider a dataset $S$ of size $n\geq 1$ and $T_{n}=T_{n}(S)$ a real-valued statistic based on it. The breakdown point for the statistic $T_{n}(S)$ is defined as $\varepsilon^{*}=p_{\infty}/n$, where $p_{\infty}=\min\{p\in\mathbb{N^{*}}:\,\sup_{S^{\prime}_{p}}|T_{n}(S)-T_{n}(S^{\prime}_{p})|=\infty\}$, the supremum being taken over all corrupted datasets $S^{\prime}_{p}$ obtained by replacing $p$ samples in $S$ by arbitrary samples. The breakdown point is a popular notion of robustness Huber and Ronchetti [2011], corresponding here to the fraction of samples that need to be corrupted to make the statistic $T_{n}$ arbitrarily large (i.e., "break down"). For example, the breakdown point of the $\alpha$-trimmed mean is given by $\lfloor\alpha n\rfloor/n$, which is approximately $\alpha$. We consider here a generalization of this notion, namely the $\tau$-breakdown point by replacing $p_{\infty}$ with $p_{\tau}=\min\{p\in\mathbb{N^{*}}\enspace\text{, }\sup_{S^{\prime}_{p}}|T_{n}(S)-T_{n}(S^{\prime}_{p})|\geq\tau\}$, where $\tau>0$ is a threshold parameter. Since our algorithm does not compute the exact $\alpha$-trimmed mean, we focus instead on determining the $\tau$-breakdown point $\varepsilon^{*}_{k}(t)$ of the partial $\alpha$-trimmed mean at iteration $t$ for each node $k$, i.e., the estimate $Z_{k}(t)$. Given that this quantity was previously shown to converge to the $\alpha$-trimmed mean, we expect that $\varepsilon^{*}_{k}(t)\leq\alpha$. The following theorem provides framing bounds for the breakdown point of the estimates of GoTrim when paired with GoRank.
@@
See Section E of the Supplementary Material for the technical proof, which relies on the idea that, when estimating a partial $\alpha$-trimmed mean, there are two sources of error: (1) the uncertainty in rank estimation, which can lead to incorrect data points being included in the mean, and (2) the delay from the gossip averaging, which requires a certain propagation time $T$ to update the network estimates. Note that, as $t\rightarrow\infty$, Theorem˜4 recovers the breakdown point of the exact $\alpha$-trimmed mean.
@@ 4.3 Numerical Experiments
Setup.
(a) Role of $\sigma_{k}^{2}/\gamma_{k}^{2}$
@@ Results.
5 Conclusion and Discussion
We introduced and analyzed two novel gossip algorithms: GoRank for rank estimation and GoTrim for trimmed mean estimation. We proved convergence rates of $\mathcal{O}(1/t)$ and established robustness guarantees for GoTrim through breakdown point analysis. Empirical results show both methods perform well on large, poorly connected networks: GoRank quickly estimates ranks, and GoTrim is robust to outliers and improves on the naive mean.
Byzantine Robustness.
In this work, we focused on robustness to data contamination in the sense of Huber’s framework. Extending these results to the more adversarial setting of Byzantine robustness remains an interesting direction. Developing a rigorous theoretical foundation for this setting is still an open problem, and we plan to address it in future work.
Asynchronous extension.
Although our analysis focuses on the synchronous setting, real-world systems are often asynchronous. We present the asynchronous version of GoRank in Appendix H. The theoretical analysis in the asynchronous setting will be carried out in an extension to this work.
Scalability.
To demonstrate the scalability of our method on large networks, we repeated the experiments from Fig. (c) in Sections 3.2 and 4.3, originally conducted with $n=500$, on larger networks with $n=1000$ and $n=5000$. The detailed results are provided in Appendix I.
Robustness to Network Disruptions.
While robustness to data contamination is important, the robustness of our proposed algorithms to network disruptions (e.g., edge or node failures, network partitioning) is equally crucial in real-world applications. In Appendix J, we provide a detailed analysis of how our current framework can be extended.
Performance on Sparse Graphs.
An interesting question is whether our algorithms perform well on sparse graphs. In practice, however, performance depends more on the graph’s connectivity than on its sparsity. To illustrate this, we present experiments in Appendix K on sparse graphs with varying levels of connectivity.
Rank-based Statistics.
GoTrim naturally extends to the decentralized estimation of rank-based statistics. This includes rank statistics [Hájek and Šidák, 1999], which are key tools in data analysis—particularly for robust hypothesis testing—as well as L-statistics (such as the Winsorized mean).
Further Discussion.
In appendix L, we provide extended discussion on several topics, including the optimality of the bounds, faster gossip algorithms, extension to multivariate data, and potential applications like robust decentralized optimization.
@@ Appendix A Gossip Algorithms for Ranking
Here, we detail two gossip algorithms for ranking, which can be considered natural competitors to GoRank. The algorithm presented first was proposed in Chiuso et al. [2011] and, to our knowledge, is the only decentralized ranking algorithm documented in the literature. The algorithm presented next can be seen as a variant of the latter, incorporating the more efficient communication scheme of GoRank.
@@
Chiuso et al. propose a gossip algorithm for distributed ranking in a general (connected) network Chiuso et al. [2011]. They demonstrate that this algorithm solves the ranking problem almost surely in finite time. However, they do not provide any convergence rate or non-asymptotic convergence results. The algorithm is outlined in Algorithm˜3 and proceeds as follows. At each time step $t$, an edge $(i,j)$ is selected, and the algorithm operates in three phases:
@@ B.1 Properties of Gossip Matrices
    The eigenvalue $\lambda_{2}(\alpha)$ satisfies $0\leq\lambda_{2}(\alpha)<1$ and $\lambda_{2}(\alpha)=1-\frac{\lambda_{2}}{\alpha|E|}$ where $\lambda_{2}$ the spectral gap (or second smallest eigenvalue) of the Laplacian.
@@
The proofs of (a), (b), (d), and (e) are omitted for brevity. For more details, we refer the reader to Boyd et al. [2006], Colin et al. [2015]. The proof of (c) follows from the fact that
@@
The proof of (f) can be found in Colin et al. [2015] (see Lemmas 1, 2 and 3). ∎
@@ B.2 Convergence Analysis of Gossip for Averaging
Algorithm 5 Gossip algorithm for estimating the standard average Boyd et al. [2006]
@@
Let us assume that $\mathcal{G}=(V,E)$ is connected and non bipartite. Then, for $\mathbf{Z}(t)$ defined in Algorithm˜5, we have that for all $k\in[n]$ :
@@
The proof is kept brief, as this is a standard result. For more details, we refer the reader to Boyd et al. [2006], Colin et al. [2015]. To prove this lemma, we will primarily use points (e) and (f) from ˜Lemma.
@@
Setting $c_{2}=1-\lambda_{2}(2)>0$ finishes the proof, as $0<\lambda_{2}(2)<1$. ∎
@@
Let us assume that $\mathcal{G}=(V,E)$ is connected and non bipartite. Then, for $\mathbf{Z}(t)$ defined in Algorithm˜5, we have that for any $t>0$ :
@@
The original proof can be found in Boyd et al. [2006].
@@
Setting $c_{2}=1-\lambda_{2}(2)>0$ finishes the proof, as $0<\lambda_{2}(2)<1$. ∎